A rectangular swimming pool measures 25ft by 50ft and has a 6 foot wide sidewalk around it. How much fencing is needed to enclose the sidewalk and what is the area of the sidewalk?
Answer by graphmatics(170)

We form a new rectangle by adding 6 feet on each end of the width and length, that is the new rectangle is 25+12=37 and the new length is 50+12=62. The perimeter of the new rectangle, the length of the sidewalk fence, is 2*37+2*62=198 feet. The area of the sidewalk is the area of the new rectangle 37*62=2294 minus the area of the pool 25*50=1250. So 2294-1250=1044.